OpenStudy (jgriffin):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

j = t + 10 t - 3 = 2j solve for j

OpenStudy (anonymous):

oops j - 3 = 2t

OpenStudy (anonymous):

oops again j-3 = 2(t-3)

OpenStudy (anonymous):

I'll try to be clearer without providing the exact answer. We need to convert the words to equations. Let j = Coach Johnson's current age. Let t = Tim's current age. The first statement says that j = t + 10. The second statement says that j-3 = 2(t-3) Solve by substitution. Substitute t + 10 for j in the second equation. t + 10 - 3 = 2(t - 3) Now solve for t.